OpenStudy (anonymous):

help me step by step? lim as x approaches 8 of (sqrt(x+1)-3)/(x-8)

OpenStudy (anonymous):

I know I can't factor, so what do I do now? Rationalize?

zepdrix (zepdrix):

\[\large \lim_{x \rightarrow 8} \frac{\sqrt{x+1}-3}{x-8}\]Hmmmm so right now we're getting an indeterminate form... So we need to find a way to cancel something out. Yesss! The way we'll do that, is by multiplying the top and bottom by the CONJUGATE of the top. Do you know what that will look like? :D

OpenStudy (anonymous):

\[\frac{ \sqrt{x+1} -3}{ x-8}*\frac{ \sqrt{x+1} +3}{ \sqrt{x+1} +3} ?\]

zepdrix (zepdrix):

Yes, good good. Let's simply the top, but DON'T multiply out the bottom, just leave the 2 terms separate for now.

OpenStudy (anonymous):

doesn't the top cancel out?

OpenStudy (anonymous):

\[\frac{ 1 }{ \sqrt{x+1}+3*x-8}\] ?

zepdrix (zepdrix):

Multiplying conjugates will give you the - Difference of Squares. I'm not sure why you think they would cancel out :D heh

OpenStudy (anonymous):

err.. sorry ;/

zepdrix (zepdrix):

\[\large (a-b)(a+b)=a^2-b^2\]

OpenStudy (anonymous):

\[\sqrt{x+1}^2-3^2?\]

zepdrix (zepdrix):

Yes c:

OpenStudy (anonymous):

\[\frac{ \sqrt{x+1}^2-3^2 }{ \sqrt{x+1}+3*x-8 }\]

zepdrix (zepdrix):

Use parentheses on the bottom :) It looks too sloppy that way. And simplify the sqrt term, see how you have a sqrt root being squared?

OpenStudy (anonymous):

\[\frac{ x-8 }{ (\sqrt{x+1}+3)(x-8) }= \sqrt{x+1}+3\]

zepdrix (zepdrix):

Yay looks good! Now it's in a nice form where we can take the limit! Since we were able to cancel some garbage out.

OpenStudy (anonymous):

=6?

zepdrix (zepdrix):

Sounds about right :D yay team.

OpenStudy (anonymous):

it says 1/6 as the answer? did I mess up somewhere

zepdrix (zepdrix):

oh lol sorry i wasn't paying attention :)

zepdrix (zepdrix):

When you made that nice cancellation. How did the thing in the denominator pop up to the top?? :O

OpenStudy (anonymous):

when I cancel it needs to become 1/.... ?

zepdrix (zepdrix):

Yah :3

zepdrix (zepdrix):

x-8 / x-8 = 1

OpenStudy (anonymous):

gotcha ;D thanks!!